What is the range of cot inverse(2x-x^2) ? (Inverse trigonometry class 12)

Grade:11

1 Answers

Ravi malvia
6 years ago
-(1/0)\small < -x2+2x \small \leq -D/4a
-(1/0)\small <-x2+2x \small \leq -4/(-4)
-(1/0)\small < -x2+2x\small \leq 1
cot-1(-1/0)\small < cot-1( -x2+2x) \small \leq cot-1(1)
 0 \small < cot-1( -x2+2x) \small \leq (22/7)÷4
so cot-1(-x2+2x ) \small \in [0,\tiny \Pi/4]